OpenStudy (anonymous):

Mix A- 2 cups concentrate and 3 cups cold water. Mix B- 1 cup concentrate and 4 cups cold water. Mix C- 4 cups concentrate and 8 cups cold water. Mix D- 3 cups concentrate and 5 cups cold water. ques.1. Which recipe will make the juice that is the most "Orangey"? Explain your answer. 2. Which will make the least orangey? again explain your answer. Need help people!

OpenStudy (anonymous):

a=2/5 b=1/5 c=1/3 d=3/8 and therefore ans is A

OpenStudy (anonymous):

I need an explanation :S

OpenStudy (anonymous):

NUM=concentrate and denominator=concentrate +cold water ihope u got it!!!!

OpenStudy (anonymous):

Nope!

OpenStudy (anonymous):

i was telling abt the fraction